### v2

@@ 1 INTRODUCTION
Vsible light communication (VLC) has attracted significant interest as a key enabling technology for future sixth generation (6G) wireless networks. The rapidly growing demand for high speed wireless connectivity and the increasing congestion and licensing cost of the radio frequency (RF) spectrum motivate the search for alternative access technologies, and VLC addresses this demand by exploiting the large optical band between approximately 400 THz and 790 THz [haas2016lifi]. In addition to this abundant bandwidth, VLC offers several attractive features, including licence free operation, high energy efficiency through light emitting diode (LED) luminaires, low electromagnetic interference, and the potential for seamless integration with existing solid state lighting infrastructure [elgala2011indoor, karunatilaka2015ledvlcsurvey]. These properties make VLC particularly well suited to high density indoor environments such as offices, hospitals, and industrial facilities, where reliable, secure, and fast throughput connectivity is required.
@@
Despite these advantages, security remains a critical concern for VLC systems. The optical signal is largely confined within the line-of-sight (LoS) region and does not penetrate opaque walls, which provides a higher degree of spatial confinement compared with conventional RF links. However, in typical indoor deployments, any user located within the illumination footprint of the LED luminaires can potentially intercept the transmitted signal, which exposes the system to unauthorised eavesdropping [Arafa_2019, mostafa2015physicallayer]. Traditional security mechanisms that rely on upper layer cryptographic protocols are often not sufficient in this context, especially in dense networks with a large number of low power internet of things (IoT), since key distribution, key renewal, and authentication procedures introduce non-negligible overhead and latency [arfaoui2020vlc_pls_survey]. These limitations have motivated the use of physical layer security (PLS) techniques, which exploit the randomness and spatial characteristics of the wireless channel to provide information theoretic security by ensuring that the achievable rate at the legitimate receiver, Bob, is strictly higher than that at the eavesdropper, Eve [bloch2011wireless, sezgin2009secrecy]. The associated performance metric, known as the secrecy capacity, quantifies the maximum rate at which perfectly secure communication can be achieved.
@@
Building upon the need for secure indoor communication, several works have investigated PLS techniques tailored to indoor VLC systems. In these studies, the inherent directionality of the optical beam and the geometric constraints of the room are exploited to enhance secrecy performance. Beamforming and null steering at the LED, often combined with artificial noise injection, have been proposed to improve the secrecy rate of multiple input single output wiretap channels in indoor scenarios [mostafa2015physicallayer, pham2024AN_VLC]. In addition, optical elements such as lenses or multibeam structures have been employed to further confine the useful signal in space and to reduce information leakage towards potential eavesdroppers [cirkinoğlu2021lens_pls, ding2022multibeam_pls]. Collectively, these approaches show that indoor VLC links can provide strong secrecy guarantees when the transmitter configuration and room geometry are carefully designed. However, the propagation environment is typically treated as fixed, and the opportunity to reconfigure the channel through additional programmable surfaces remains largely unexplored.
@@
Recently, intelligent reflecting surface (IRS) technology has been introduced into VLC systems as a means to actively manipulate the indoor propagation environment and to overcome the limitations of fixed geometry. An IRS typically consists of a planar array of passive or low power reflecting elements whose optical reflection properties can be individually controlled, which allows the incident light to be redirected in a programmable manner [aboagye2021irs_vlc, sun2022irs_resource]. In indoor VLC, such programmable surfaces can be deployed on walls or ceilings to create additional controllable paths between the transmitter and the users. From a PLS perspective, this provides a new degree of freedom, since appropriate configuration of the surface can concentrate the received power at the legitimate user while suppressing the signal at the eavesdropper location, which leads to a significant improvement in secrecy performance [abumarshoud2023icc_pls, qian2024secureVLC_irs].
@@
However, most existing works on IRS-assisted VLC security rely on simplified channel models. A common assumption is that the composite channel formed by the LoS link and the IRS reflected links is frequency flat, or that the temporal dispersion introduced by the reflected paths can be neglected [aboagye2021irs_vlc, sun2022irs_resource, qian2024secureVLC_irs]. This assumption is reasonable for low rate narrowband modulation, yet it becomes inaccurate in practical high speed VLC systems that employ wideband schemes such as direct current optical orthogonal frequency division multiplexing (DCO-OFDM). In such systems, the path length differences between the direct LoS component and the multiple non-line-of-sight (NLoS) components reflected by the IRS introduce non-negligible time delays, which give rise to frequency selective fading and inter-symbol interference (ISI). These effects can degrade performance if they are not properly taken into account in the system and security analysis. A few recent channel modelling studies have started to characterise the impact of IRS-induced time delay on the frequency domain behaviour of VLC links [chen2023freqdomainIRS, HAIDER2025102703].
@@ 1.1 Related Work
PLS for indoor VLC was first studied mainly under the assumption that eavesdroppers act independently. Early works considered single cell scenarios in which one LED serves a single user in the presence of randomly located non-colluding eavesdroppers. In this setting, secrecy capacity and secrecy outage expressions were derived, and the impact of key parameters such as the transmit optical power, the receiver field of view, and the user locations was characterised [pan2017secureVLC_random, cho2018secure_beamforming]. These results showed that indoor VLC can achieve strong secrecy performance thanks to the spatial confinement of the optical beam and the possibility of shaping the coverage region through power control and beamforming. Subsequent studies extended the analysis to more realistic LiFi networks with multiple access points and user mobility, and evaluated secrecy under imperfect channel state information and device orientation, which provided a more accurate picture of practical deployments [abumarshoud2021realistic_secrecy]. A variety of transmit techniques were proposed, including multiple input single output beamforming and artificial noise, power allocation strategies, and shaped optical beams, all aimed at improving the secrecy rate when eavesdroppers are non-colluding and experience independent channels [arfaoui2020vlc_pls_survey, cho2018secure_beamforming].
The assumption that eavesdroppers do not cooperate may, however, be optimistic in dense indoor environments where multiple malicious users can share their observations. This has motivated a line of work on PLS with colluding eavesdroppers. In the context of VLC, [cho2018colluding_vlc] analysed the secrecy performance of a downlink system with randomly located colluding eavesdroppers that combine their received signals, and derived closed form expressions for the secrecy outage probability using tools from stochastic geometry, while [abumarshoud2021realistic_secrecy] considered a multi cell LiFi network and showed that colluding eavesdroppers can significantly reduce the achievable secrecy capacity, especially when realistic receiver orientation is taken into account. More broadly, the detrimental effect of collusion on secrecy capacity has been quantified in the wireless literature, where information theoretic analyses demonstrated that cooperation among eavesdroppers shrinks the security region and tightens secrecy constraints [sezgin2009secrecy, bloch2011wireless]. Overall, these studies highlighted that indoor VLC is vulnerable to colluding eavesdroppers, and that system design based solely on non-colluding models may overestimate the available secrecy capacity.
@@
irs have also been investigated in the context of PLS. By properly controlling the reflection properties of a large number of low cost elements, an IRS can strengthen the legitimate channel. Motivated by this concept, several works have introduced IRS concepts into indoor VLC. Programmable optical reflectors mounted on walls or ceilings have been used to create additional controllable paths between the transmitter and the receivers, and have been shown to improve coverage and enhance secrecy, for example in non-orthogonal multiple access (NOMA) based VLC systems and in downlink LiFi scenarios with non-colluding eavesdroppers [abumarshoud2023icc_noma_irs_vlc, qian2024secureVLC_irs, abumarshoud2024irs_vlc_secure, HAIDER2025102703]. In most of these designs, the composite channel formed by the LoS link and the IRS reflected links was modelled as frequency flat, and the propagation delay associated with the reflected paths was not explicitly taken into account.
In practical wideband VLC systems that employ intensity modulation with direct detection, the different path lengths between the light source, the IRS elements, and the receivers introduce non-negligible time delays, which lead to frequency selective fading and ISI. Our recent work [leveraging2024irs_delay_pls] was the first to exploit this IRS induced time delay as a degree of freedom for enhancing PLS in indoor VLC. In that work, a single eavesdropper was considered, and an optimisation problem was formulated to allocate the IRS elements in order to maximise the secrecy capacity by creating constructive ISI at the legitimate user and destructive ISI at the eavesdropper. The numerical results showed that a time delay aware use of the IRS can provide substantial secrecy gains compared to configurations that ignore the time delay. However, the analysis in [leveraging2024irs_delay_pls] was limited to one eavesdropper, and did not address scenarios with multiple non-colluding and colluding eavesdroppers under a unified time delay IRS framework. To the best of our knowledge, a comparative secrecy capacity analysis for colluding and non-colluding eavesdroppers in time delay IRS-assisted VLC systems has not been reported so far.
@@ 1.2 Motivation and Contributions
In this paper, we use the IRS induced time delay to improve the PLS of indoor VLC system. Instead of neglecting the time delay in VLC channels, we shape the channel impulse response so that the reflected signals add constructively at Bob and destructively at the eavesdroppers. By choosing how the IRS elements are allocated, strong reflected paths align with the LoS component at Bob, while they become misaligned at the eavesdroppers, which creates harmful ISI for them. This effect is particularly important when the eavesdroppers are closer to the transmitter than Bob. We also consider a general threat model with $K$ eavesdroppers that may act independently or collude by combining their received signals. The secrecy capacity maximisation problem that follows from this idea involves discrete allocation of the IRS elements and is non-convex and combinatorial. To address this, we use a reinforcement learning (RL)method based on proximal policy optimisation (PPO), which learns an allocation strategy for the IRS elements from interaction with the environment.
The main contributions of this paper are as follows:
  • We formulate a secrecy capacity maximisation problem for an indoor VLC system assisted by IRS, where the discrete allocation of the IRS elements shapes the channel frequency response to create constructive interference for Bob and destructive interference for non-colluding and colluding eavesdroppers.
  • We apply a PPO based RL algorithm with suitable state, action, and reward definitions, so that the agent learns the IRS allocation strategy that increases the long term secrecy performance.
  • We present simulation results that illustrate the secrecy capacity behaviour of the proposed IRS-induced time-delay for different numbers and locations of eavesdroppers, and show that positive secrecy can be maintained even in worst case geometric configurations where the colluding eavesdroppers have stronger channels than Bob.
@@ 2 System Model
As shown in Fig. 1, an indoor VLC scenario is considered. A single LED transmitter, denoted as Alice, is mounted at the centre of the ceiling and provides both illumination and high speed data transmission. The confidential message is intended for a single legitimate user, Bob, who is located inside the room and also receives illumination from the same LED. Due to the widebeam of the LED light, any receiver within the illumination footprint can in principle intercept the transmitted signal. A security threat is therefore modelled in the form of $K$ passive eavesdroppers, collected in the set $\mathcal{E}=\{E_{1},\dots,E_{K}\}$. These eavesdroppers are assumed to be arbitrarily located within the room, to receive the illumination, and to remain silent, so that their presence is not detected by Alice.
To enhance the PLS of the legitimate link, an IRS is mounted on one of the walls. The IRS consists of $N_{\text{irs}}$ passive reflecting elements with separation distance $D$ between adjacent elements. The incident optical signal from the LED illuminates the IRS, and each element reflects part of this signal towards the interior of the room. In this manner, additional NLoS path is created between Alice and each receiver, with propagation distances that depend on the positions of the IRS elements and of the users. The different path lengths associated with these links give rise to different arrival times at the receivers and contribute to the overall time dispersion of the channel. In the remainder of this section, the end-to-end optical channel for each receiver is expressed as the superposition of the direct LoS component and the IRS-assisted NLoS components, and the corresponding channel impulse response (CIR) and channel frequency response (CFR) expressions used in the secrecy capacity analysis are derived.
@@ 2.1 LoS Channel
where $A$ is the physical photodetector (PD) area and $d_{\scriptscriptstyle{k,\text{LED}}}$ is the LoS distance from Alice to user $k$. The model parameters include the LED irradiance angle $\phi_{\scriptscriptstyle{k,\text{LED}}}$, the receiver incidence angle $\psi_{\scriptscriptstyle{k,\text{LED}}}$, the optical filter gain $\mathbb{G}_{\scriptscriptstyle f}$, and the optical concentrator gain $\mathbb{G}_{\scriptscriptstyle c}$. The channel gain depends strongly on the geometry. The term $\cos^{m}(\phi)$ models the LED radiation pattern, while $\cos(\psi)$ accounts for the receiver orientation. The field of view constraint $\Psi_{\scriptscriptstyle\text{FoV}}$ imposes a hard cut off, since any ray arriving at an angle greater than $\Psi_{\scriptscriptstyle\text{FoV}}$ is rejected and leads to zero gain. The concentrator gain $\mathbb{G}_{\scriptscriptstyle c}$ is a function of the lens refractive index $\xi$, and $\Psi_{\scriptscriptstyle\text{FoV}}$ is expressed as
$\mathbb{G}_{c}=\begin{cases}\frac{\xi^{2}}{\sin^{2}\!\bigl(\Psi_{\scriptscriptstyle\text{FoV}}\bigr)}\,,&0\leq\psi_{\scriptscriptstyle{k,\text{LED}}}\leq\Psi_{\scriptscriptstyle\text{FoV}}\,,\\ 0\,,&\text{otherwise}\end{cases}$ (2)

### v3

@@ 1 INTRODUCTION
Vsible light communication (VLC) has attracted significant interest as a key enabling technology for future sixth generation (6G) wireless networks. The rapidly growing demand for high speed wireless connectivity and the increasing congestion and licensing cost of the radio frequency (RF) spectrum motivate the search for alternative access technologies, and VLC addresses this demand by exploiting the large optical band between approximately 400 THz and 790 THz [16]. In addition to this abundant bandwidth, VLC offers several attractive features, including licence free operation, high energy efficiency through light emitting diode (LED) luminaires, low electromagnetic interference, and the potential for seamless integration with existing solid state lighting infrastructure [15, 19]. These properties make VLC particularly well suited to high density indoor environments such as offices, hospitals, and industrial facilities, where reliable, secure, and fast throughput connectivity is required.
@@
Despite these advantages, security remains a critical concern for VLC systems. The optical signal is largely confined within the line-of-sight (LoS) region and does not penetrate opaque walls, which provides a higher degree of spatial confinement compared with conventional RF links. However, in typical indoor deployments, any user located within the illumination footprint of the LED luminaires can potentially intercept the transmitted signal, which exposes the system to unauthorised eavesdropping [7, 20]. Traditional security mechanisms that rely on upper layer cryptographic protocols are often not sufficient in this context, especially in dense networks with a large number of low power internet of things (IoT), since key distribution, key renewal, and authentication procedures introduce non-negligible overhead and latency [8]. These limitations have motivated the use of physical layer security (PLS) techniques, which exploit the randomness and spatial characteristics of the wireless channel to provide information theoretic security by ensuring that the achievable rate at the legitimate receiver, Bob, is strictly higher than that at the eavesdropper, Eve [9, 26]. The associated performance metric, known as the secrecy capacity, quantifies the maximum rate at which perfectly secure communication can be achieved.
@@
Building upon the need for secure indoor communication, several works have investigated PLS techniques tailored to indoor VLC systems. In these studies, the inherent directionality of the optical beam and the geometric constraints of the room are exploited to enhance secrecy performance. Beamforming and null steering at the LED, often combined with artificial noise injection, have been proposed to improve the secrecy rate of multiple input single output wiretap channels in indoor scenarios [20, 22]. In addition, optical elements such as lenses or multibeam structures have been employed to further confine the useful signal in space and to reduce information leakage towards potential eavesdroppers [cirkinoğlu2021lens_pls, 14]. Collectively, these approaches show that indoor VLC links can provide strong secrecy guarantees when the transmitter configuration and room geometry are carefully designed. However, the propagation environment is typically treated as fixed, and the opportunity to reconfigure the channel through additional programmable surfaces remains largely unexplored.
@@
Recently, intelligent reflecting surface (IRS) technology has been introduced into VLC systems as a means to actively manipulate the indoor propagation environment and to overcome the limitations of fixed geometry. An IRS typically consists of a planar array of passive or low power reflecting elements whose optical reflection properties can be individually controlled, which allows the incident light to be redirected in a programmable manner [1, 25]. In indoor VLC, such programmable surfaces can be deployed on walls or ceilings to create additional controllable paths between the transmitter and the users. From a PLS perspective, this provides a new degree of freedom, since appropriate configuration of the surface can concentrate the received power at the legitimate user while suppressing the signal at the eavesdropper location, which leads to a significant improvement in secrecy performance [4, 23].
@@
However, most existing works on IRS-assisted VLC security rely on simplified channel models. A common assumption is that the composite channel formed by the LoS link and the IRS reflected links is frequency flat, or that the temporal dispersion introduced by the reflected paths can be neglected [1, 25, 23]. This assumption is reasonable for low rate narrowband modulation, yet it becomes inaccurate in practical high speed VLC systems that employ wideband schemes such as direct current optical orthogonal frequency division multiplexing (DCO-OFDM). In such systems, the path length differences between the direct LoS component and the multiple non-line-of-sight (NLoS) components reflected by the IRS introduce non-negligible time delays, which give rise to frequency selective fading and inter-symbol interference (ISI). These effects can degrade performance if they are not properly taken into account in the system and security analysis. A few recent channel modelling studies have started to characterise the impact of IRS-induced time delay on the frequency domain behaviour of VLC links [11, 17].
@@ 1.1 Related Work
PLS for indoor VLC was first studied mainly under the assumption that eavesdroppers act independently. Early works considered single cell scenarios in which one LED serves a single user in the presence of randomly located non-colluding eavesdroppers. In this setting, secrecy capacity and secrecy outage expressions were derived, and the impact of key parameters such as the transmit optical power, the receiver field of view (FoV), and the user locations was characterised [21, 13]. These results showed that indoor VLC can achieve strong secrecy performance thanks to the spatial confinement of the optical beam and the possibility of shaping the coverage region through power control and beamforming. Subsequent studies extended the analysis to more realistic LiFi networks with multiple access points and user mobility, and evaluated secrecy under imperfect channel state information (CSI) and device orientation, which provided a more accurate picture of practical deployments [6]. A variety of transmit techniques were proposed, including multiple input single output beamforming and artificial noise, power allocation strategies, and shaped optical beams, all aimed at improving the secrecy rate when eavesdroppers are non-colluding and experience independent channels [8, 13].
The assumption that eavesdroppers do not cooperate may, however, be optimistic in dense indoor environments where multiple malicious users can share their observations. This has motivated a line of work on PLS with colluding eavesdroppers. In the context of VLC, [12] analysed the secrecy performance of a downlink system with randomly located colluding eavesdroppers that combine their received signals, and derived closed form expressions for the secrecy outage probability using tools from stochastic geometry, while [6] considered a multi cell LiFi network and showed that colluding eavesdroppers can significantly reduce the achievable secrecy capacity, especially when realistic receiver orientation is taken into account. More broadly, the detrimental effect of collusion on secrecy capacity has been quantified in the wireless literature, where information theoretic analyses demonstrated that cooperation among eavesdroppers shrinks the security region and tightens secrecy constraints [26, 9]. Overall, these studies highlighted that indoor VLC is vulnerable to colluding eavesdroppers, and that system design based solely on non-colluding models may overestimate the available secrecy capacity.
@@
irs have also been investigated in the context of PLS. By properly controlling the reflection properties of a large number of low cost elements, an IRS can strengthen the legitimate channel. Motivated by this concept, several works have introduced IRS concepts into indoor VLC. Programmable optical reflectors mounted on walls or ceilings have been used to create additional controllable paths between the transmitter and the receivers, and have been shown to improve coverage and enhance secrecy, for example in non-orthogonal multiple access (NOMA) based VLC systems and in downlink LiFi scenarios with non-colluding eavesdroppers [5, 23, 3, 17]. In most of these designs, the composite channel formed by the LoS link and the IRS reflected links was modelled as frequency flat, and the propagation delay associated with the reflected paths was not explicitly taken into account.
In practical wideband VLC systems that employ intensity modulation with direct detection, the different path lengths between the light source, the IRS elements, and the receivers introduce non-negligible time delays, which lead to frequency selective fading and ISI. Our recent work [18] was the first to exploit this IRS induced time delay as a degree of freedom for enhancing PLS in indoor VLC. In that work, a single eavesdropper was considered, and an optimisation problem was formulated to allocate the IRS elements in order to maximise the secrecy capacity by creating constructive ISI at the legitimate user and destructive ISI at the eavesdropper. The numerical results showed that a time delay aware use of the IRS can provide substantial secrecy gains compared to configurations that ignore the time delay. However, the analysis in [18] was limited to one eavesdropper, and did not address scenarios with multiple non-colluding and colluding eavesdroppers under a unified time delay IRS framework. To the best of our knowledge, a comparative secrecy capacity analysis for colluding and non-colluding eavesdroppers in time delay IRS-assisted VLC systems has not been reported so far.
@@ 1.2 Motivation and Contributions
In this paper, we use the IRS induced time delay to improve the PLS of an indoor VLC system. Rather than neglecting the time delay in VLC channels, we shape the channel impulse response so that the reflected signals add constructively at Bob and destructively at the eavesdroppers. By carefully choosing how the IRS elements are allocated, strong reflected paths align with the LoS component at Bob, while becoming misaligned at the eavesdroppers, creating harmful ISI that degrades their reception. This effect is particularly pronounced when the eavesdroppers are located closer to the transmitter than Bob, making it a powerful tool in challenging geometric configurations. Furthermore, to strengthen the threat model, we consider a general scenario with $K$ eavesdroppers that may act independently or collude by combining their received signals. However, the secrecy capacity maximisation problem that arises from this formulation involves discrete allocation of the IRS elements, rendering it non-convex and combinatorial in nature. To overcome this challenge, we employ a reinforcement learning (RL) method based on proximal policy optimisation (PPO), which learns an effective allocation strategy for the IRS elements through direct interaction with the environment. The main contributions of this paper are as follows:
  • We formulate a secrecy capacity maximisation problem for an indoor VLC system assisted by an IRS, where the discrete allocation of the IRS elements shapes the channel frequency response to create constructive interference for Bob and destructive interference for both non-colluding and colluding eavesdroppers.
  • Moreover, we apply a PPO based RL algorithm with carefully designed state, action, and reward definitions, enabling the agent to learn an IRS allocation strategy that consistently improves long-term secrecy performance.
  • Finally, we present simulation results that illustrate the secrecy capacity behaviour of the proposed IRS-induced time-delay approach for varying numbers and locations of eavesdroppers, demonstrating that positive secrecy can be maintained even in worst-case geometric configurations where colluding eavesdroppers possess stronger channels than Bob.
@@ 2 System Model
As shown in Fig. 1, an indoor VLC scenario is considered. In this environment, a single LED transmitter, Alice, is mounted at the centre of the ceiling to provide simultaneous illumination and high-speed data transmission. The confidential message is intended exclusively for a single legitimate user, Bob; however, the widebeam nature of the LED light ensures that the signal covers a large footprint, allowing any receiver within this area to potentially intercept the transmission. Consequently, a security threat is modelled in the form of $K$ passive eavesdroppers, denoted by the set $\mathcal{E}=\{E_{1},\dots,E_{K}\}$, who are arbitrarily located within the room. These eavesdroppers are assumed to remain physically silent to avoid detection by the transmitter. Although these eavesdroppers do not reveal their presence through uplink transmissions, we assume that their CSI is perfectly known at the transmitter to investigate the fundamental secrecy limits, which follows the established analytical convention in PLS research to evaluate performance under a theoretical worst-case security benchmark. This assumption is supported by recent advancements in optical wireless channel estimation, such as the closed-form location and orientation estimation framework presented in [10], which demonstrates the practical feasibility of acquiring such detailed CSI in VLC systems.
To enhance the PLS of the legitimate link, an IRS is mounted on one of the walls. The IRS consists of $N_{\text{irs}}$ passive reflecting elements with separation distance $D$ between adjacent elements. The incident optical signal from the LED illuminates the IRS, and each element reflects part of this signal towards the interior of the room by adjusting its orientation. In particular, each element is oriented to steer its reflection toward a selected receiver direction, hence a single element is assumed to be allocated to one user at a time, in line with mirror array based optical IRS operation [24]. In this manner, additional NLoS path is created between Alice and each receiver, with propagation distances that depend on the positions of the IRS elements and of the users. The different path lengths associated with these links give rise to different arrival times at the receivers and contribute to the overall time dispersion of the channel. In the remainder of this section, the end-to-end optical channel for each receiver is expressed as the superposition of the direct LoS component and the IRS-assisted NLoS components, and the corresponding channel impulse response (CIR) and channel frequency response (CFR) expressions used in the secrecy capacity analysis are derived.
@@ 2.1 LoS Channel
where $A$ is the physical photodetector (PD) area and $d_{\scriptscriptstyle{k,\text{LED}}}$ is the LoS distance from Alice to user $k$. The model parameters include the LED irradiance angle $\phi_{\scriptscriptstyle{k,\text{LED}}}$, the receiver incidence angle $\psi_{\scriptscriptstyle{k,\text{LED}}}$, the optical filter gain $\mathbb{G}_{\scriptscriptstyle f}$, and the optical concentrator gain $\mathbb{G}_{\scriptscriptstyle c}$. The channel gain depends strongly on the geometry. The term $\cos^{m}(\phi)$ models the LED radiation pattern, while $\cos(\psi)$ accounts for the receiver orientation. The FoV constraint $\Psi_{\scriptscriptstyle\text{FoV}}$ imposes a hard cut off, since any ray arriving at an angle greater than $\Psi_{\scriptscriptstyle\text{FoV}}$ is rejected and leads to zero gain. The concentrator gain $\mathbb{G}_{\scriptscriptstyle c}$ is a function of the lens refractive index $\xi$, and $\Psi_{\scriptscriptstyle\text{FoV}}$ is expressed as
$\mathbb{G}_{c}=\begin{cases}\frac{\xi^{2}}{\sin^{2}\!\bigl(\Psi_{\scriptscriptstyle\text{FoV}}\bigr)}\,,&0\leq\psi_{\scriptscriptstyle{k,\text{LED}}}\leq\Psi_{\scriptscriptstyle\text{FoV}}\,,\\ 0\,,&\text{otherwise},\end{cases}$ (2)